Problem 4 Sketch the sinusoid x(t)-2 cos(2n(10)t + π/2) over the range-0.1 0.2.
Answer #1

The above function will be a cosine graph of the form,

r(t) = Acos(2nft o)

so, the amplitude of the above waveform is 2.0 units, the frequency of the wave is 10 Hz and the phase difference is quarter of a wave (pi/2). The time period of the wave is T = 1/f = 0.1s.

The graph looks like:

0/05 0.05 015 0,25

Here, the horizontal axis is given in radians and the shaded region is the region -0.1 < t < 0.2.